# DevBench: A Comprehensive Benchmark for Software Development

### v2

@@ Appendix C Discussions
C.1 Model Capacity
@@
The detailed results broken down on different languages are illustrated in Figure 5. Makefiles and Gradle files are pivotal in C/C++ and Java repositories, orchestrating file organization, compilation, linking, and managing dependencies. However, models struggle in accurately generating these files. Smaller models often misconstrue the syntax and structure, leading to irrelevant content like misplaced Linux commands or repeating given prompt instructions. Although larger models, such as GPT-4-Turbo, show improvements, they still fall short, exemplified by frequent “missing separator” errors in Makefiles. Similarly, Gradle files often lack crucial dependency declarations, a critical oversight leading to Java compilation failures. Another frequent Gradle problem is omitting essential tasks. Although given with necessary “tests” and “acceptanceTests” prompts, the model often misses these in the Gradle file, but can correct them upon review.
@@ C.4 Limitations in Testing
Appendix D Repositories statistics in DevBench
@@
Table 8: Repositories statistics in DevBench

### v3

@@ Appendix D Software Design Evaluation
Appendix E Experimental Discussions
E.1 Model Capacity
@@
LLMs often face challenges in generating accurate Makefile for C/C++ and Gradle files for Java. Frequently, the generated files are deficient in critical components like source code files, necessary dependencies and essential tasks. In C/C++ and Java repositories, approximately 90% of compilation and execution errors can be attributed to these issues. We find that even GPT-4-Turbo occasionally fail in basic syntax errors on compilation files. This is potentially caused by insufficient training data related to these compilation tools.
@@ E.4 Limitations in Testing
To prevent meaningless but executable testing code, structured test templates with explicit instructions and incomplete assertion statements can potentially guide and force models toward meaningful test generation.
Appendix F Repositories statistics in DevEval
Table 8 shows the repository statistics within DevEval.
@@ Appendix F Repositories statistics in DevEval
Table 8: Repository statistics within DevEval